Let { (1,1,1,1) , (1,2,1,2) } be a linearly independent sbuset of a vector space V4. Extend it to a basic of V4
1
Expert's answer
2021-08-26T18:22:57-0400

It is enough to find a matrix "L=\left(\begin{array}{llll}1&1&1&1\\1&2&1&2\\x_1&x_2&x_3&x_4\\y_1&y_2&y_3&y_4\end{array}\right)" wit nonzero determinant. Set "x_1=0", "x_2=0", "x_3=1", "x_4=0"; "y_1=0", "y_2=0", "y_3=0" , "y_4=1". The determinant of the matrix is: "det(L)=\det(A)\det(B)", where "A=\left(\begin{array}{ll}1&1\\1&2\end{array}\right)" and "B=\left(\begin{array}{ll}1&0\\0&1\end{array}\right)". "det(A)=1", "det(B)=1". "det(L)=1". Thus, the basis is: "(1\,\,1\,\,1\,\,1)", "(1\,\,2\,\,1\,\,2)", "(0\,\,0\,\,1\,\,0)", "(0\,\,0\,\,0\,\,1)".